# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

### v1

@@ Appendix C Linear Separability Analysis
Table 11: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
@@
Figure 7: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers. The linear separability gap measures how much a non-linear transformation improves classifier performance compared to a simple linear mapping. For inflection prediction, the gap is consistently minimal (mostly within ±$\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network. By contrast, the linear separability gap for lemma prediction is relatively large (0.1–0.3 percentage points) and positive across most models
Figure 8: Intrinsic dimensionality curves for all 10 models. Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers. The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model. Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
Figure 9: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps). The full model is 928k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Early checkpoints exhibit higher lexeme accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
Figure 10: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps). The full model is 143k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Lexeme accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.

### v4

@@ I.3 Cross-Linguistic Patterns
Table 11: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model using Multi-Layer Perceptron (MLP) classifiers (English). Lemmas are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale. MLP classifiers provide consistent improvements over linear regression across all POS categories, though the frequency-dependent performance patterns persist.
Model Linear Regression MLP
Positive Base Plural Singular Positive Base Plural Singular
(n=300) (n=2,074) (n=3) (n=3,947) (n=300) (n=2,074) (n=3) (n=3,947)
mT5-Base 0.739 0.913 0.436 0.962 0.783 0.919 0.231 0.961
Qwen2.5-1.5B 0.785 0.929 0.034 0.969 0.801 0.924 0.092 0.967
Qwen2.5-1.5B-Instruct 0.779 0.925 0.034 0.964 0.803 0.923 0.057 0.967
Qwen2.5-7B 0.824 0.937 0.310 0.970 0.828 0.929 0.310 0.969
Qwen2.5-7B-Instruct 0.819 0.936 0.299 0.970 0.823 0.928 0.276 0.969
Goldfish Chinese 0.793 0.912 0.000 0.958 0.816 0.915 0.000 0.957
Table 12: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Chinese). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Preposition Other
(n=1,179) (n=564) (n=108) (n=22) (n=20) (n=50)
mT5-Base 0.838 0.828 0.786 0.762 0.920 0.726
Qwen2.5-1.5B 0.810 0.797 0.746 0.715 0.872 0.699
Qwen2.5-1.5B-Instruct 0.813 0.799 0.748 0.713 0.873 0.700
Qwen2.5-7B 0.887 0.882 0.846 0.847 0.915 0.817
Qwen2.5-7B-Instruct 0.886 0.877 0.843 0.835 0.913 0.811
Goldfish Chinese 0.883 0.878 0.845 0.875 0.954 0.858
Table 13: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression classifiers (Chinese). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Preposition Other
(n=1,179) (n=564) (n=108) (n=22) (n=20) (n=50)
mT5-Base 0.698 0.712 0.564 0.571 0.884 0.569
Qwen2.5-1.5B 0.748 0.761 0.658 0.668 0.826 0.669
Qwen2.5-1.5B-Instruct 0.735 0.745 0.643 0.643 0.814 0.655
Qwen2.5-7B 0.815 0.826 0.749 0.745 0.848 0.750
Qwen2.5-7B-Instruct 0.815 0.822 0.747 0.734 0.845 0.744
Goldfish Chinese 0.766 0.771 0.647 0.621 0.912 0.682
Table 14: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers (Chinese). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Superlative Singular Comparative
(n=417) (n=517) (n=1,720) (n=839) (n=1,076) (n=52) (n=3,197) (n=141)
mT5-Base 0.908 0.941 0.940 0.960 0.882 0.572 0.962 0.636
Qwen2.5-1.5B 0.849 0.889 0.922 0.914 0.888 0.657 0.953 0.796
Qwen2.5-1.5B-Instruct 0.844 0.887 0.922 0.910 0.889 0.659 0.952 0.795
Qwen2.5-7B 0.892 0.922 0.939 0.947 0.909 0.826 0.962 0.878
Qwen2.5-7B-Instruct 0.915 0.934 0.945 0.962 0.924 0.866 0.968 0.909
Goldfish German 0.938 0.941 0.955 0.979 0.916 0.542 0.968 0.708
Table 15: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (German). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Superlative Singular Comparative
(n=417) (n=517) (n=1,720) (n=839) (n=1,076) (n=52) (n=3,197) (n=141)
mT5-Base 0.921 0.945 0.948 0.959 0.884 0.723 0.967 0.770
Qwen2.5-1.5B 0.890 0.915 0.930 0.940 0.897 0.831 0.958 0.892
Qwen2.5-1.5B-Instruct 0.888 0.914 0.930 0.938 0.898 0.825 0.957 0.897
Qwen2.5-7B 0.912 0.932 0.944 0.956 0.913 0.868 0.964 0.924
Qwen2.5-7B-Instruct 0.925 0.941 0.950 0.966 0.928 0.901 0.970 0.936
Goldfish German 0.947 0.957 0.964 0.978 0.923 0.817 0.970 0.896
Table 16: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (German). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Linear Regression MLP
Noun Verb Adjective Other Noun Verb Adjective Other
(n=1,262) (n=395) (n=406) (n=12) (n=1,262) (n=395) (n=406) (n=12)
mT5-Base 0.685 0.662 0.568 0.750 0.611 0.602 0.486 0.723
Qwen2.5-1.5B 0.743 0.725 0.715 0.775 0.721 0.700 0.687 0.711
Qwen2.5-1.5B-Instruct 0.740 0.722 0.715 0.766 0.722 0.698 0.687 0.704
Qwen2.5-7B 0.821 0.809 0.808 0.829 0.795 0.786 0.783 0.814
Qwen2.5-7B-Instruct 0.815 0.803 0.803 0.821 0.795 0.785 0.782 0.813
Goldfish German 0.720 0.747 0.701 0.769 0.758 0.772 0.742 0.769
Table 17: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (German). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Singular
(n=688) (n=776) (n=1,833) (n=857) (n=1,457) (n=5,169)
mT5-Base 0.934 0.912 0.879 0.908 0.954 0.970
Qwen2.5-1.5B 0.933 0.858 0.896 0.903 0.958 0.967
Qwen2.5-1.5B-Instruct 0.930 0.852 0.893 0.898 0.958 0.966
Qwen2.5-7B 0.955 0.918 0.918 0.931 0.965 0.975
Qwen2.5-7B-Instruct 0.951 0.913 0.915 0.928 0.964 0.974
Goldfish French 0.942 0.955 0.937 0.930 0.968 0.976
Table 18: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (French). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Singular
(n=688) (n=776) (n=1,833) (n=857) (n=1,457) (n=5,169)
mT5-Base 0.957 0.937 0.911 0.935 0.957 0.977
Qwen2.5-1.5B 0.954 0.905 0.914 0.925 0.965 0.968
Qwen2.5-1.5B-Instruct 0.954 0.902 0.911 0.924 0.965 0.968
Qwen2.5-7B 0.966 0.936 0.930 0.937 0.970 0.976
Qwen2.5-7B-Instruct 0.962 0.931 0.926 0.934 0.970 0.975
Goldfish French 0.974 0.967 0.945 0.942 0.973 0.979
Table 19: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (French). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Linear Regression MLP
Noun Verb Adjective Other Noun Verb Adjective Other
(n=1,496) (n=406) (n=358) (n=15) (n=1,496) (n=406) (n=358) (n=15)
mT5-Base 0.708 0.577 0.605 0.799 0.755 0.560 0.636 0.820
Qwen2.5-1.5B 0.754 0.725 0.673 0.824 0.807 0.765 0.751 0.853
Qwen2.5-1.5B-Instruct 0.750 0.718 0.671 0.820 0.824 0.776 0.768 0.869
Qwen2.5-7B 0.840 0.814 0.764 0.869 0.856 0.825 0.794 0.884
Qwen2.5-7B-Instruct 0.833 0.805 0.758 0.860 0.851 0.818 0.792 0.883
Goldfish French 0.749 0.758 0.661 0.811 0.894 0.869 0.813 0.888
Table 20: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (French). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Superlative Singular Comparative
(n=690) (n=456) (n=1,192) (n=455) (n=1,333) (n=3) (n=3,316) (n=23)
mT5-Base 0.930 0.978 0.975 0.957 0.877 0.000 0.977 0.799
Qwen2.5-1.5B 0.925 0.946 0.974 0.938 0.923 0.015 0.966 0.835
Qwen2.5-1.5B-Instruct 0.924 0.943 0.974 0.934 0.921 0.015 0.966 0.817
Qwen2.5-7B 0.949 0.966 0.979 0.958 0.948 0.094 0.977 0.872
Qwen2.5-7B-Instruct 0.951 0.974 0.980 0.970 0.948 0.080 0.980 0.918
Goldfish Russian 0.940 0.950 0.976 0.931 0.921 0.000 0.976 0.867
Table 21: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (Russian). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Superlative Singular Comparative
(n=690) (n=456) (n=1,192) (n=455) (n=1,333) (n=3) (n=3,316) (n=23)
mT5-Base 0.959 0.978 0.969 0.966 0.904 0.000 0.978 0.849
Qwen2.5-1.5B 0.952 0.955 0.972 0.948 0.933 0.089 0.970 0.899
Qwen2.5-1.5B-Instruct 0.950 0.954 0.973 0.947 0.933 0.089 0.969 0.911
Qwen2.5-7B 0.963 0.964 0.978 0.960 0.951 0.246 0.979 0.910
Qwen2.5-7B-Instruct 0.961 0.970 0.978 0.966 0.949 0.126 0.980 0.924
Goldfish Russian 0.965 0.972 0.978 0.948 0.943 0.000 0.977 0.934
Table 22: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (Russian). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Linear Regression MLP
Noun Verb Adjective Other Noun Verb Adjective Other
(n=982) (n=333) (n=275) (n=4) (n=982) (n=333) (n=275) (n=4)
mT5-Base 0.660 0.614 0.542 0.648 0.492 0.484 0.387 0.426
Qwen2.5-1.5B 0.777 0.712 0.759 0.720 0.712 0.696 0.716 0.647
Qwen2.5-1.5B-Instruct 0.772 0.704 0.756 0.720 0.710 0.689 0.717 0.643
Qwen2.5-7B 0.854 0.790 0.843 0.812 0.798 0.794 0.813 0.749
Qwen2.5-7B-Instruct 0.845 0.778 0.835 0.807 0.794 0.785 0.809 0.744
Goldfish Russian 0.795 0.723 0.764 0.676 0.810 0.776 0.759 0.657
Table 23: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Russian). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Singular
(n=154) (n=51) (n=401) (n=168) (n=33) (n=632)
mT5-Base 0.860 0.911 0.928 0.966 0.837 0.952
Qwen2.5-1.5B 0.808 0.802 0.721 0.928 0.861 0.892
Qwen2.5-1.5B-Instruct 0.809 0.817 0.720 0.941 0.878 0.899
Qwen2.5-7B 0.865 0.879 0.810 0.966 0.903 0.909
Qwen2.5-7B-Instruct 0.850 0.874 0.796 0.960 0.886 0.900
Goldfish Turkish 0.847 0.915 0.880 0.964 0.872 0.963
Table 24: Breakdown of inflection classification accuracy for each model by inflection type using Linear Regression classifiers (Turkish). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Base 3rd person Positive Past Plural Singular
(n=154) (n=51) (n=401) (n=168) (n=33) (n=632)
mT5-Base 0.755 0.760 0.848 0.922 0.515 0.949
Qwen2.5-1.5B 0.770 0.767 0.667 0.919 0.765 0.914
Qwen2.5-1.5B-Instruct 0.762 0.757 0.662 0.917 0.766 0.913
Qwen2.5-7B 0.853 0.845 0.791 0.956 0.875 0.937
Qwen2.5-7B-Instruct 0.845 0.844 0.786 0.956 0.875 0.932
Goldfish Turkish 0.832 0.879 0.870 0.957 0.834 0.957
Table 25: Breakdown of inflection classification accuracy for each model by inflection type using Multi-Layer Perceptron (MLP) classifiers (Turkish). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
Model Linear Regression MLP
Noun Verb Adjective Other Noun Verb Adjective Other
(n=221) (n=53) (n=104) (n=13) (n=221) (n=53) (n=104) (n=13)
mT5-Base 0.866 0.823 0.921 0.955 0.215 0.421 0.374 0.637
Qwen2.5-1.5B 0.834 0.805 0.866 0.877 0.307 0.439 0.449 0.693
Qwen2.5-1.5B-Instruct 0.816 0.791 0.860 0.874 0.305 0.439 0.448 0.691
Qwen2.5-7B 0.871 0.850 0.900 0.904 0.595 0.625 0.695 0.809
Qwen2.5-7B-Instruct 0.850 0.823 0.883 0.885 0.579 0.613 0.678 0.800
Goldfish Turkish 0.929 0.904 0.940 0.969 0.386 0.550 0.477 0.808
Table 26: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model, using Linear Regression and Multi-Layer Perceptron (MLP) classifiers (Turkish). Lemmas are grouped by their POS tags (e.g., , Noun, Verb, Adjective). Accuracies are calculated over all examples for a given group across all layers. Counts (n) are derived from a single representative layer for each group. All accuracy values are on a 0–1 scale.
@@
Figure 22: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers for English. The linear separability gap measures how much a non-linear transformation improves classifier performance compared to a simple linear mapping. For inflection prediction, the gap is consistently minimal (mostly within $\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network. By contrast, the linear separability gap for lemma prediction is relatively large (0.1–0.3 percentage points) and positive across most models
Cross-Linguistic Linear Separability Gap
(a) Chinese - Inflection
(b) Chinese - Lemma
(c) German - Inflection
(d) German - Lemma
(e) French - Inflection
(f) French - Lemma
(g) Russian - Inflection
(h) Russian - Lemma
(i) Turkish - Inflection
(j) Turkish - Lemma
Figure 23: Cross-linguistic linear separability gap showing performance advantage of MLP classifiers over linear classifiers across model layers for five additional languages. For inflectional features, mT5 and Goldfish models show slight positive gaps (indicating modest benefits from non-linear classification), while Qwen2.5 variants show slight negative gaps (indicating linear classifiers are sufficient or superior). For lexical features, all models show negative gaps that are most pronounced in early layers, suggesting that linear regression with regularization consistently outperforms MLPs for lexical classification across all model families and languages.
Figure 24: Lemma (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps) for English. The full model is 928k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Early checkpoints exhibit higher lemma accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
Figure 25: Lemma (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps) for English. The full model is 143k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Lemma accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.